A rectangular playground is 88 meters wide. It is 33 times as long as it is wide.What is the area of the playground?

Answer by Alan3354(69443)

A rectangular playground is 88 meters wide. It is 33 times as long as it is wide.What is the area of the playground?
---------
W = 88
L = 88*33 = 2704 meters
----
Area = L*W = 237952 sq meters